OpenStudy (donnie1999):

HELP PLEASE!!! Write the expression in factored form 100m^4-n^2 @aloud

OpenStudy (anonymous):

=100m4−n2

OpenStudy (anonymous):

im not sure if its right tho

OpenStudy (jack1):

\(\Large 100m^4-n^2\) is this ur equation? or is it: \(\Large 100m^{(4-n^2)}\)??

OpenStudy (donnie1999):

the first one is the equation @Jack1

OpenStudy (anonymous):

100m^4−n^2

OpenStudy (jack1):

\(\Large 100m^4-n^2\) so reverse FOIL this then u get our ax^2 + bx + c ==> 100m^4-n^2 so a = 100, b = 0 , c = n turns into a factored form of ( x+something ) (x - something )

OpenStudy (jack1):

100m4−n2 a = 100 b = 0 c = -n ( x+n ) (x - n ) n's are solved to equal C now for a 100 = 10 x 10 so (10x + n ) (10x - n ) proof ==> FOIL it out First 10x * 10x = 100x^2 Outers 10x * (-n) = -10xn Inners 10x * (n) = 10xn Last n * (-n) = -n^2 add all those together 100m4−n2